# UL-UNAS: Ultra-Lightweight U-Nets for Real-Time Speech Enhancement via Network Architecture Search

### v1

@@ 2.5 Neural architecture search for speech enhancement
While NAS methods have outperformed manually designed architectures on many tasks such as image classification, object detection, and semantic segmentation, there are few works on applying NAS in SE. NAS-TasNet (Lee et al., 2022) investigates applying both gradient descent and RL-based NAS methods on a classic speech separation model, Conv-TasNet (Luo and Mesgarani, 2019), and achieves higher performance with fewer parameters. However, the resulting model remains too large to be considered as lightweight. In Vafeiadis et al. (2024), a quantization-aware NAS method is used to design an ultra-low memory SE system with just 9.2k parameters. However, the search space is built upon a trivial convolutional recurrent encoder-decoder architecture family, which limits its performance upper bound and hinders its potential to achieve more advanced capabilities. In this paper, we apply NAS to discover the optimal configuration for constructing an ultra-lightweight U-Net using our elaborately designed blocks.
3 Methodology
3.1 Problem formulation
@@
where $X(t,f)$, $S(t,f)$, and $N(t,f)$ denote the complex spectrograms of the noisy speech, the clean speech, and the noise, respectively, with time index $t$ and frequency index $f$. The masking method is employed to obtain the enhanced speech, as we have found it performs better than mapping when the model size is restricted. Although estimating cIRM is widely adopted in many SE models to recover the phase, we have observed that it is prone to yielding an estimated mask with nearly zero imaginary parts, meaning that it is downgraded to IRM. This is in alignment with the phenomena elucidated in Yin et al. (2020). We attribute this to the insufficient capacity of the model. Constrained by the complexity of the model, many methods with greater performance potentials have proven to be inapplicable. Therefore, we resort to a straightforward yet effective IRM-based masking method to recover the clean spectrogram from the noisy one, which can be expressed as:
$\tilde{S}(t,f)=X(t,f)\odot M(t,f),$ (2)
where $\odot$ denotes element-wise multiplication and $\tilde{S}(t,f)$ represents the enhanced spectrogram.
3.2 Model architecture
We extract the key concepts from various efficient convolutional block designs discussed in Sec. 2.2 to construct our basic blocks and evaluate their effectiveness in the field of SE. After identifying the most promising designs, we augment them with our proposed two novel boosting components: APReLU and cTFA. The details of these modules will be presented in the following sections.
@@ 3.2 Model architecture
3.2.1 Basic blocks
We explore various efficient convolutional block designs, including:
  • 1.
@@ 3.2.1 Basic blocks
  • 2.
    DWS block: The depthwise separable block from MobileNet (Howard, 2017), which decomposes a standard convolution to a depthwise and a pointwise convolution.
  • 3.
    Ghost block: The lightweight block introduced in GhostNet (Han et al., 2020), which uses a cost-effective depthwise convolution to create ghost features;
  • 4.
    Rep block: The efficient block introduced in MobileOne (Vasu et al., 2023), which applies reparameterization techniques on the DWS block to improve performance without extra computation during inference;
  • 5.
    MB block: The inverted residual block from MobileNetV2 (Sandler et al., 2018), incorporating an additional pointwise convolution and residual connection based on the DWS block;
  • 6.
    Star block: The optimal variant of star blocks explored in StarNet (Ma et al., 2024), leveraging the star operation for implicit high-dimensional representative capacity.
@@
We classify these blocks into three categories based on the number of convolutional layers. Within each category, the most promising block is selected for augmentation with our proposed APReLU and cTFA modules, and subsequently serves as a candidate for NAS.
3.2.2 Affine PReLU
The APReLU activation function is a combination of an affine transformation and a PReLU activation, formulated as:
@@ 3.2.2 Affine PReLU
where $x\in\mathbb{R}^{C\times T\times F}$ represents the input feature, with $C,T,F$ denoting the channel, time, and frequency axis lengths, respectively. $\gamma,\beta\in\mathbb{R}^{C\times F}$ is the frequency-dependent scaling factor and bias of the affine transformation, and $\alpha\in\mathbb{R}^{C}$ is the slope-controlling parameter in PReLU. The $\gamma$ coefficient is initialized to all ones, and the $\beta$ coefficient is initialized to zeros so that the affine transformation starts as an identity mapping. The $\alpha$ coefficient is initialized to 0.25. All three coefficients are learnable.
Example APReLU functions are illustrated in Fig. 2. It should be noted that, in contrast to conventional PReLU, our proposed APReLU enables a wider range of activation modes. Importantly, the scaling factor and bias of the affine transformation are frequency-dependent, enabling activating different modes for different frequency bands.
3.2.3 Causal time-frequency attention
TFA (Zhang et al., 2022) is a lightweight attention module for salient T-F speech distribution modeling using two parallel branches, namely time attention and frequency attention. Concerning an intermediate T-F representation, time-frame-wise and frequency-wise statistical features are respectively generated through average pooling over the frequency and time dimensions. Subsequently, stacked one-dimensional (1D) convolutional layers are employed to model the distribution and produce the attention map for each statistical feature. A final two-dimensional (2D) T-F attention map is obtained via a tensor multiplication operation. Although it demonstrates effectiveness in being flexibly integrated with existing backbone networks to improve performance, the average pooling along the time dimension gives rise to non-causality, thereby hampering its application in real-time SE models. In TRA (Rong et al., 2024), the non-causal branch is discarded so that the final attention map is frequency-independent, which unfortunately leads to a certain degree of performance deterioration.
@@ 3.2.3 Causal time-frequency attention
Our proposed cTFA module, as shown in Fig. 3, consists of two parallel branches: frequency-independent temporal attention (FITA) and frequency-dependent temporal attention (FDTA). The FITA branch is the same as the TRA module, which consists of a GRU layer, a fully connected (FC) layer, and a Sigmoid activation. The FDTA branch contains two stacked 2D convolutional layers with a PReLU activation inserted in between, followed by a Sigmoid activation. The first convolutional layer expands the channel dimension by a factor of $r$, while the second layer restores it to its original size. In this paper, the kernel sizes of the two convolutions are set to (3,1), and the expansion factor $r$ is set to 5.
Given $V\in\mathbb{R}^{C\times T\times F}$ as the input feature, the frequency-independent temporal attention map $A_{T}\in\mathbb{R}^{C\times T}$ is first computed as follows:
$Z_{T}$ $=\frac{1}{F}\sum_{f=1}^{F}V^{2}(c,t,f),$ (5)
$A_{T}$ $=\text{FITA}(Z_{T}).$ (6)
On the other hand, the frequency-dependent temporal attention map $A_{F}\in\mathbb{R}^{1\times T\times F}$ is computed as follows:
$Z_{F}$ $=\frac{1}{C}\sum_{c=1}^{C}V^{2}(c,t,f),$ (7)
$A_{F}$ $=\text{FDTA}(Z_{F}).$ (8)
The final multi-dimensional attention map $A_{TF}\in\mathbb{R}^{C\times T\times F}$ is computed as:
$A_{TF}(c,t,f)=A_{T}(c,t)\times A_{F}(t,f).$ (9)
@@
$\tilde{V}=V\odot A_{TF},$ (10)
where $\odot$ denotes element-wise multiplication operation.
3.2.4 Model prototype
We adopt the GTCRN framework to construct our model prototype, as shown in Fig. 4. We preserve the band merging (BM), band splitting (BS), and grouped dual-path RNN (G-DPRNN) modules, while replacing the original convolutional blocks with the aforementioned basic blocks. Note that for blocks in the decoder, the convolutional layers are replaced with transposed variants accordingly. After identifying the three most promising block candidates, the two advantageous components are integrated to form three extended convolutional blocks: all activations are substituted with APReLU, and a cTFA module is added at the end. Examples are shown in Fig. 5.
Figure 4: Overall architecture of the model prototype, with blank blocks representing the building blocks to be equipped.
3.3 Neural architecture search
We optimize the network configuration within a search space composed of the aforementioned extended blocks using NAS. To enhance search efficiency, we design a relatively small search space with respect to the encoder. Once the encoder is determined, the entire network can be ascertained. The encoder is partitioned into a sequence of our pre-defined extended blocks, with each block spanning a per-block sub-search space. Specifically, the sub-search space for block $i$ comprises 5 configurable nodes, each offering multiple options:
  • 1.
@@ 3.3 Neural architecture search
The stride size of a block refers to the stride along the frequency dimension in the standard or depthwise convolution within the block. The group number of a block refers to the number of groups in the standard or pointwise convolution within the block. When the group number is 2, a channel shuffle (Zhang et al., 2018) operation is applied afterward. The encoder architecture comprises five blocks, generating 25 searchable nodes and creating a search space of approximately $10^{12}$ possible configurations. We employ the same RL-based search algorithm proposed in Tan et al. (2019). However, in contrast to optimizing for latency, we focus on optimizing MACS since we do not target any specific hardware device.
Figure 6: Diagram of MACS-aware neural architecture search.
@@
where $\omega_{+}$ and $\omega_{-}$ are application-specific constants. Eventually, the parameters of the controller are updated by maximizing the expected reward using proximal policy optimization (Schulman et al., 2017). The aforementioned procedure is repeated until the parameters of the controller converge.
3.4 Loss function
@@
$\begin{split}\mathcal{L}=\alpha\mathcal{L}_{\text{SISNR}}(\tilde{s},s)+(1-\beta)\mathcal{L}_{\text{mag}}(\tilde{S},S)\\ +\beta\left(\mathcal{L}_{\text{real}}(\tilde{S},S)+\mathcal{L}_{\text{imag}}(\tilde{S},S)\right),\end{split}$ (13)
where $\tilde{s}$ and $s$ are the enhanced and clean waveforms. $\tilde{S}$ and $S$ are the enhanced and clean spectrograms, respectively. $\alpha$ and $\beta$ are set to 0.01 and 0.3 respectively. Each term in the aforementioned formula is calculated as follows:

### v2

@@ II-E Neural architecture search for speech enhancement
While NAS methods have outperformed manually designed architectures on many tasks such as image classification, object detection, and semantic segmentation, there are few works on applying NAS in SE. NAS-TasNet [24] investigates applying both gradient descent and RL-based NAS methods on a classic speech separation model, Conv-TasNet [27], and achieves higher performance with fewer parameters. However, the resulting model remains too large to be considered lightweight. In [52], a quantization-aware NAS method is used to design an ultra-low memory SE system with just 9.2k parameters. However, the search space is defined by a trivial U-Net architecture, which limits its performance upper bound and hinders its potential to achieve more advanced capabilities. In this paper, we apply NAS to discover the optimal configuration for constructing an ultra-lightweight U-Net, where the search space is defined by our elaborately designed, high-performance blocks.
III Methodology
Figure 1: Overall architecture of UL-UNAS.
III-A Problem formulation
@@
where $X(t,f)$, $S(t,f)$, and $N(t,f)$ denote the complex spectrograms of the noisy speech, the clean speech, and the noise, respectively, with time index $t$ and frequency index $f$. We employ the masking method to obtain the enhanced speech, as it has been shown to perform better than mapping when the model size is limited. Although estimating cIRM is widely adopted in many SE models to recover the phase, we have observed that it is prone to yielding an estimated mask with nearly zero imaginary parts, meaning that it is downgraded to IRM. This is in alignment with the phenomena elucidated in [63]. We attribute this to the insufficient capacity of the ultra-lightweight models. Therefore, we resort to a straightforward yet effective IRM-based masking method to recover the clean spectrogram from the noisy one, which can be expressed as:
$\tilde{S}(t,f)=X(t,f)\otimes M(t,f),$ (2)
where $\otimes$ denotes element-wise multiplication and $\tilde{S}(t,f)$ represents the enhanced spectrogram.
III-B Model Overview
We adopt the GTCRN framework as the model backbone, as illustrated in Fig. 1. We retain the band merging (BM), band splitting (BS), and grouped dual-path RNN (G-DPRNN) modules, while replacing the original convolutional blocks with our elaborately designed blocks: XConv, XDWS, and XMB blocks, which are described in the following sections. For blocks in the decoder, convolutional layers are replaced with transposed variants accordingly. The number of blocks, their arrangement, and the associated hyperparameter configurations are all determined via NAS. The model takes the noisy log-power spectrogram as input and predicts a magnitude mask, constrained within the range of (0,1), to recover the enhanced magnitude. The enhanced spectrogram is then obtained by combining this magnitude with the original noisy phase.
III-C Model Components
In this section, we first investigate the most efficient convolutional block designs for SE, as surveyed in Sec. II-B. We then introduce two novel components proposed in this work: APReLU and cTFA, which are designed to augment model representational capacity. The final encoder and decoder architectures are constructed by integrating these components into the selected base blocks.
@@ III-C Model Components
III-C1 Base blocks
We explore various efficient convolutional block designs as follows.
@@ III-C1 Base blocks
  • DWS block: The depthwise separable block from MobileNet [15], which decomposes a standard convolution to a depthwise and a pointwise convolution.
  • Ghost block: The lightweight block introduced in GhostNet [10], which uses a cost-effective depthwise convolution to create ghost features;
  • Rep block: The efficient block introduced in MobileOne [56], which applies reparameterization techniques on the DWS block to improve performance without extra computation during inference;
  • MB block: The inverted residual block from MobileNetV2 [43], incorporating an additional pointwise convolution and residual connection based on the DWS block;
  • Star block: The optimal variant of star blocks explored in StarNet [29], leveraging the star operation for implicit high-dimensional representative capacity.
@@
We classify these blocks into three categories based on the number of convolutional layers inside. Within each category, the most promising block is selected for augmentation with our proposed APReLU and cTFA modules, and subsequently serves as a candidate for NAS. Based on our ablation studies in Sec. V-A1, the Conv, DWS, and MB blocks were selected, from which we constructed the XConv, XDWS, and XMB blocks, respectively, as illustrated in Fig. 3.
III-C2 Affine PReLU
As speech characteristics and noise distributions vary substantially across frequency bands, we argue that a frequency-dependent activation function is more appropriate than conventional frequency-agnostic alternatives for speech enhancement. To this end, we introduce the APReLU activation function, which combines a frequency-dependent affine transformation with a PReLU activation, formulated as:
@@ III-C2 Affine PReLU
where $x\in\mathbb{R}^{C\times T\times F}$ represents the input feature, with $C,T,F$ denoting the channel, time, and frequency axis lengths, respectively. $\gamma,\beta\in\mathbb{R}^{C\times F}$ is the frequency-dependent scaling factor and bias of the affine transformation, and $\alpha\in\mathbb{R}^{C}$ is the slope-controlling parameter in PReLU. The $\gamma$ coefficient is initialized to all ones and the $\beta$ coefficient to zeros, so that the affine transformation starts as an identity mapping, which allows the module to start from a residual connection and gradually learn adaptations. The $\alpha$ coefficient is initialized to 0.25 following the commonly used default in PyTorch, providing a stable and widely adopted starting point. All three coefficients are learnable.
Example APReLU functions are illustrated in Fig. 4. As compared to conventional PReLU, our proposed APReLU supports a broader range of activation modes. Notably, the frequency-dependent scaling factors and biases enable the activation to recalibrate the energy across different frequency bands, thereby offering the potential to more effectively enhance speech components while suppressing noise.
@@
III-C3 Causal time-frequency attention
TFA [65] is a lightweight attention module for salient T-F speech distribution modeling using two parallel branches, namely time attention (TA) and frequency attention (FA). Concerning an intermediate T-F representation, time-frame-wise and frequency-wise statistical features are respectively generated through average pooling over the frequency and time dimensions. Subsequently, stacked one-dimensional (1D) convolutional layers are employed to model the distribution and produce the attention map for each statistical feature. A final two-dimensional (2D) T-F attention map is obtained via a tensor multiplication operation. Although it demonstrates effectiveness in being flexibly integrated with existing backbone networks to improve performance, the average pooling along the time dimension gives rise to non-causality, thereby hampering its application in real-time SE models. In TRA [41], the non-causal branch is discarded so that the final attention map is frequency-independent, which unfortunately leads to a certain degree of performance deterioration. However, we argue that the essential role of FA lies in modeling frequency dependencies, and this goal does not inherently require time-pooling. Accordingly, our cTFA module specifically replaces the time-pooling-based FA branch with a channel-pooling-based one. This design provides an effective alternative for frequency modeling while inherently ensuring causality.
Our proposed cTFA module, built upon TRA, is designed to further capture frequency dependencies and generate a finer-grained attention map. To circumvent the non-causality issue in TFA, we avoid any time-pooling operations and instead introduce a lightweight frequency-attention mechanism based on channel-pooling features. As shown in Fig. 5, cTFA consists of two parallel branches: TA and FA. The TA branch is the same as the TRA module, which consists of a GRU layer, a fully connected (FC) layer, and a Sigmoid activation. The FA branch adopts an analogous architecture but replaces the GRU with a bidirectional GRU (BiGRU) to achieve stronger spectral modeling capability. To balance representational capacity and computational cost, we introduce an efficient frequency folding operation. This process transforms the channel-pooled feature map (of size $B\times T\times F$) into a compact representation (of size $BT\times H\times R$) by reshaping the frequency dimension. Specifically, the $F$ frequency bins are consolidated into $H$ frequency groups, each encompassing a local spectral context of size $R$ (fixed to 4 in this work, with $F=H\cdot R$). When $F$ is not divisible by $R$, zero-padding is applied along the frequency dimension. This folding operation allows the subsequent BiGRU to model long-range dependencies along the $H$ axis and local spectral context along the $R$ axis more effectively and efficiently.
Given $V\in\mathbb{R}^{B\times C\times T\times F}$ as the input feature, the time attention map $A_{T}\in\mathbb{R}^{B\times C\times T}$ is first computed as follows:
$Z_{T}$ $=\frac{1}{F}\sum_{f=1}^{F}V^{2}(b,c,t,f),$ (5)
$A_{T}$ $=\text{TA}(Z_{T}).$ (6)
On the other hand, the frequency attention map $A_{F}\in\mathbb{R}^{B\times T\times F}$ is computed as follows:
$Z_{F}$ $=\frac{1}{C}\sum_{c=1}^{C}V^{2}(b,c,t,f),$ (7)
$A_{F}$ $=\text{FA}(Z_{F}).$ (8)
The final multi-dimensional attention map $A_{TF}\in\mathbb{R}^{B\times C\times T\times F}$ is computed as:
$A_{TF}(b,c,t,f)=A_{T}(b,c,t)\cdot A_{F}(b,t,f).$ (9)
@@ III-C3 Causal time-frequency attention
$\tilde{V}=V\otimes A_{TF},$ (10)
where $\otimes$ denotes the element-wise multiplication operation.
III-D Neural architecture search
Figure 6: Diagram of MACS-aware neural architecture search.
We optimize the network configuration within a search space composed of the aforementioned extended blocks using NAS. To improve search efficiency, we design a relatively small search space with respect to the encoder. Once the encoder is determined, the entire network can be ascertained. The encoder is partitioned into a sequence of our pre-defined extended blocks, each spanning a per-block sub-search space. Specifically, the sub-search space for block $i$ comprises 5 configurable nodes, each offering multiple options:
@@ III-D Neural architecture search
The stride size of a block refers to the stride along the frequency dimension in the standard or depthwise convolution within the block. The group number of a block refers to the number of groups in the standard or pointwise convolution within the block. When the group number is 2, a channel shuffle [67] operation is applied afterward. The encoder architecture comprises five blocks, generating 25 searchable nodes and creating a search space of approximately $10^{12}$ possible configurations. We employ the same RL-based search algorithm proposed in [50]. However, in contrast to optimizing for latency, we focus on optimizing MACS since we do not target any specific hardware device.
@@
where $\omega_{+}$ and $\omega_{-}$ are application-specific constants. Eventually, the parameters of the controller are updated by maximizing the expected reward using proximal policy optimization [45]. The aforementioned procedure is repeated until the parameters of the controller converge.
III-E Loss function
@@
$\begin{split}\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{SISNR}}(\tilde{s},s)+(1-\lambda_{2})\mathcal{L}_{\text{mag}}(\tilde{S},S)\\ +\lambda_{2}\left(\mathcal{L}_{\text{real}}(\tilde{S},S)+\mathcal{L}_{\text{imag}}(\tilde{S},S)\right),\end{split}$ (13)
where $\tilde{s}$ and $s$ are the enhanced and clean waveforms. $\tilde{S}$ and $S$ are the enhanced and clean spectrograms, respectively. $\lambda_{1}$ and $\lambda_{2}$ are set to 0.01 and 0.3, respectively, as these values were empirically determined to balance the contributions of each loss component and were also validated in GTCRN [41]. Each term in the aforementioned formula is calculated as follows:
@@
IV Experiments
IV-A Datasets
We evaluate our proposed model using two datasets. The first is the large-scale DNS3 dataset [38], which includes a diverse collection of clean speech, noise, and RIRs. Additionally, we incorporate the Mandarin corpus from DiDiSpeech [8]. During data mixing, clean speech is convolved with a randomly selected RIR and then mixed with randomly chosen noise clips under an SNR range of -5 to $15\text{\,}\mathrm{d}\mathrm{B}$. The training target is obtained by preserving the first 100 ms of reflections. A total of 720,000 pairs of 10-second noisy-clean data are created for training, while 1,000 pairs are generated for validation and another 1,000 pairs for testing, respectively. For the ablation study, a smaller subset ($10\%$ of the complete dataset, with each utterance cropped to 4 seconds) is sampled to facilitate faster convergence.
The second dataset is the VCTK-DEMAND dataset [53], which contains paired clean and pre-mixed noisy speech. The training and test set consists of 11,572 utterances from 28 speakers and 824 utterances from two speakers, respectively. 1,572 utterances in the training set are selected for validation. All the utterances are sampled at $16\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$.
IV-B Implementation details
IV-B1 Parameter configurations for model prototype
The BM module adopts the same setup in GTCRN, which maps the 192 high-frequency bands to 64 ERB bands, while keeping the 65 low-frequency bands unaltered, leading to a 129-dimensional compressed feature map. For the ablation study, we construct a model prototype whose encoder comprises five identical blocks, selected from the basic or extended blocks as described in Sec. III-C1. All blocks have 16 channels and a kernel size of (3,3). The first two blocks have a stride of 2 in the frequency dimension to downsample the feature map, while the remaining three blocks have a stride of 1. The decoder is a mirror version of the encoder, where all the standard or depthwise convolutional layers are replaced with transposed ones with the same parameter configurations.
IV-B2 Training configurations
The STFT is computed using a Hanning window with a length of 32 ms, a hop length of 16 ms, and an FFT size of 512 points. The log-power spectrogram is fed into the model, which outputs an estimated IRM to obtain the enhanced magnitude. The enhanced speech is then obtained by combining the enhanced magnitude with the noisy phase. The models are trained using the Adam optimizer [20] with two different schedulers. For the DNS3 dataset, a linear-warm-up cosine-annealing scheduler is used. The learning rate increases linearly from $10^{-6}$ to $10^{-3}$ over the first 25,000 steps and then decays cosine-wise until 250,000 steps, at which point the training procedure ends. A batch size of 32 is used for the full DNS3 dataset, while a batch size of 8 is used for the DNS3 subset. For the VCTK-DEMAND dataset, a performance-based learning rate scheduler is applied. The learning rate starts at $10^{-3}$ and is halved if the validation loss fails to decrease for 5 consecutive epochs. For this dataset, a batch size of 4 is used.
IV-B3 Neural architecture search configurations
The controller cell has an embedding size of 100 and an LSTM hidden size of 200. It is trained using the Adam optimizer with an initial learning rate of 0.001. The multi-processing technique is employed to sample and evaluate multiple architectures concurrently. For each episode, 40 architectures are sampled and trained on 8 RTX 3090 GPUs, with 5 models per GPU and a batch size of 8 for each model. Each model is trained on the DNS3 subset for 20 epochs, using 5,000 randomly selected noisy-clean pairs per epoch. Validation is conducted only during the final three epochs, and the average perceptual evaluation of speech quality (PESQ) [37] score is used as the $Q(m)$ in Eq. (11). The baseline score $Q_{0}$ is set to 1, and the target MACS $M_{T}$ is set to 30M [^1] The computational complexity of all models was evaluated using the ptflops toolkit, with a fixed input length of 1 second.. The weight constants $\omega_{+}$ and $\omega_{-}$ are set to -0.15 and 0, respectively. If the reward does not increase for 5 consecutive epochs, the learning rate is halved.
IV-C Baselines and evaluation metrics
We compare our model with the latest SOTA ultra-lightweight SE models, including GTCRN [41] and LiSenNet [61]. In addition, we include scaled-down versions of CRN [47] (CRN-small) and DPCRN [21] (DPCRN-small), as well as DeepFilterNet [44], to provide a more comprehensive comparison. The CRN-small and DPCRN-small models are implemented by ourselves, following the original papers’ configurations while proportionally reducing the number of channels for model compression. For DeepFilterNet, we use the official public implementation. All baselines are retrained under the same training setup as our UL-UNAS model.
When comparing on the VCTK-DEMAND dataset, some models from this benchmark are also considered, including FullSubNet [12], FastFullSubNet [11], and GaGNet [25]. We use the statistics reported in the original papers if available; otherwise, we retrain these models using the official implementations with our training configurations to ensure fairness.
We use three intrusive metrics for performance evaluation: the scale-invariant signal-to-noise ratio (SISNR) [42], the PESQ, and the extended short-time objective intelligibility (ESTOI) [19]. Additionally, two non-intrusive metrics, DNSMOS P.808 [39] and DNSMOS P.835 [40], are employed to provide a more comprehensive comparison. DNSMOS P.835 includes three sub-metrics: OVRL for overall speech quality, SIG for signal quality, and BAK for background noise quality.